a rectangular play ground has two semi circles added to its out side its smaller side as diameter if the sides of the rectangle are 120 meter and 21 meter find the area of the play ground?
Answer by nerdybill(7384)

a rectangular play ground has two semi circles added to its out side its smaller side as diameter if the sides of the rectangle are 120 meter and 21 meter find the area of the play ground?
.
"rectangular area" = 120*21 = 2520 square meters
.
"two semi circles area" = 2((1/2)(pi)(21/2)^2)
"two semi circles area" = (pi)(21/2)^2
"two semi circles area" = (3.14)(10.5)^2
"two semi circles area" = (3.14)(110.25)
"two semi circles area" = 346.185 square meters
.
"area of play ground" = "rectangular area" + "two semi circles area"
"area of play ground" = 2520 + 346.185
"area of play ground" = 2866.185 square meters
.